OpenStudy (anonymous):

Holly wants to save money for an emergency. Holly invests 700 in an account that pays an interest rate of 6.75%. How many years will it take for the account to reach $16,300?

OpenStudy (anonymous):

16300=700*1.0675^t t=(log(163)-log7)/log(1.0675)

OpenStudy (anonymous):

Thanks ;3